OpenStudy (pokemon23):

the diagonal of a rectangle is 13in . One side is 12in. Find other side.

OpenStudy (pokemon23):

a^2+b^2=c^2

OpenStudy (anonymous):

13^2 - 12^2 = 169 - 144 = 25 sqrt(25) = 5

OpenStudy (pokemon23):

12^2+b^2=13^2 right?

OpenStudy (anonymous):

yes

OpenStudy (pokemon23):

13^2 is the hypotenuse?

OpenStudy (anonymous):

13 is the hypotenuse

OpenStudy (anonymous):

5 is the other side

OpenStudy (pokemon23):